1.The smith's want to install a rectangular swimming pool. It cost $5.00 a square foot for concrete. If the pool is 70 feet across, 25 feet wide and 15 feet deep, what is the total cost to concrete the pool?

Answer by Theo(13342)

concrete is $5.00 a square foot.
presumably the width of the concrete wall is taken into account.
since it's not mentioned, i'll assume it's implied in the $5.00 per square foot figure.
concrete is poured on the sides of the pool and the bottom of the pool.
2 sides of the pool are 70 * 15 = 1050 square feet each for a total of 2100 square feet.
2 sides of the pool are 25 * 15 = 325 square feet each for a total of 750 square feet.
the bottom of the pool is 70 * 25 = 1750 square feet.
the total amount of concrete required is for all 4 sides plus the bottom of the pool which is equal to:
2100 + 850 + 1750 = 4600 square feet.